# Let’s Predict Sentence by Sentence

### v1

@@ 3.5 Results
Note that we measure computational costs across the full latent pipeline, including classifier and decoder components, which remain unoptimized.To see the cost with a lightweight classifier, please refer to Appendix D. Thus, reported efficiency gains represent conservative estimate. Across tasks, Continuous inference achieves 1.5–2.5$\times$ better efficiency compared to token-level CoT. Notably, we highlight that even Discretized inference outperform CoT in longer reasoning tasks (e.g., Blocksworld w/ average trace length $R\sim 9.1$: 52.26 GFLOPs vs. 58.69 GFLOPs). We expect this efficiency gap to grow as the length of reasoning trace increases.
@@ Appendix C Computation Complexity Analysis
Attention Complexity under KV-caching
@@
    Chain-of-Thought (CoT). Each step emits $L$ new tokens into the context. Before step $t$, the context length is $N_{0}+(t-1)L$, so
@@
    Language-Grounded Mode. Each step (i) processes only latents in the main chain $\bigl{(}\mathcal{O}(R^{2})\bigr{)}$ and (ii) decodes and re-encodes an $L$-token sentence $\bigl{(}\mathcal{O}(L^{2}R)\bigr{)}$, yielding
@@
    Pure Latent Mode. Each step adds one latent vector; attending over $t-1$ latents gives
@@
Summary of leading-order costs:
@@
In addition to attention cost, every decoded or re-encoded token incurs feed-forward (MLP) computation. More specifically:
@@
  • Language-Grounded: With a semantic encoder, each step decodes and re-encodes $L$ tokens on compact codes—processing $2L$ tokens per step for an MLP cost of $\mathcal{O}(LR)$. If instead a contextual encoder must re-attend over up to $N_{0}+(t-1)L$ tokens each pass, it incurs an additional $\mathcal{O}(LR^{2})$ MLP overhead, which can erode attention savings unless the encoder is shallow or non-autoregressive.
@@
Under KV-caching, the Language-Grounded mode—with a semantic encoder—adds an $\mathcal{O}(L^{2}R)$ decode/re-encode overhead, but makes it ideal for tasks sensitive to error-compounding or instability (i.e. Mathematics.) In contrast, the Pure Latent mode eliminates all token-level context (attention $\mathcal{O}(R^{2})$, MLP $\mathcal{O}(R)$), offering maximal efficiency when possible.
@@ Appendix F Evaluation Prompt
Stage GPT-2 Small GPT-2 Medium GPT-2 Large (LoRA) $r{=}256$, $a{=}1024$)

### v2

# Latent Reasoning via Sentence Embedding Prediction

@@ 3.5 Results
[…]
@@ Appendix C Computation Complexity Analysis
Attention Complexity under KV‐caching
@@
    Chain‐of‐Thought (CoT). Each step emits $L$ new tokens into the context. Before step $t$, the context length is $N_{0}+(t-1)L$, so
@@
    Language‐Grounded Mode. Each step (i) processes only latents in the main chain $\bigl(\mathcal{O}(R^{2})\bigr)$ and (ii) decodes and re‐encodes an $L$-token sentence $\bigl(\mathcal{O}(L^{2}R)\bigr)$, yielding
@@
    Pure Latent Mode. Each step adds one latent vector; attending over t1t-1 latents gives
@@
Summary of leading‐order costs:
@@
In addition to attention cost, every decoded or re‐encoded token incurs feed‐forward (MLP) computation. More specifically:
@@
  • Language‐Grounded: With a semantic encoder, each step decodes and re‐encodes $L$ tokens on compact codes—processing $2L$ tokens per step for an MLP cost of $\mathcal{O}(LR)$. If instead a contextual encoder must re‐attend over up to $N_{0}+(t-1)L$ tokens each pass, it incurs an additional $\mathcal{O}(LR^{2})$ MLP overhead, which can erode attention savings unless the encoder is shallow or non‐autoregressive.
@@
Under KV‐caching, the Language‐Grounded mode—with a semantic encoder—adds an $\mathcal{O}(L^{2}R)$ decode/re‐encode overhead, but makes it ideal for tasks sensitive to error‐compounding or instability (i.e. Mathematics.) In contrast, the Pure Latent mode eliminates all token‐level context (attention $\mathcal{O}(R^{2})$, MLP $\mathcal{O}(R)$), offering maximal efficiency when possible.
@@ Appendix F Evaluation Prompt
Stage GPT-2 Small GPT-2 Medium GPT-2 Large (LoRA) r=256r{=}256, a=1024a{=}1024)